OpenStudy (anonymous):

help! will give medal and fan to the first person to awnser add the following x/x+10 + 12/x3 -100x

OpenStudy (anonymous):

x^3 or 3x?

OpenStudy (anonymous):

X^3

OpenStudy (whpalmer4):

\[\frac{x}{x+10}+\frac{12}{x^3}-100x\]Is that it?

OpenStudy (anonymous):

Yeah thats the question.

OpenStudy (whpalmer4):

If so, you need to make a common denominator. This one is a bit trickier than the one we did earlier, because one of the terms has no denominator whatsoever, and there are three terms, not two. No matter :-) We'll make the third term have a denominator of 1: \[\frac{x}{x+10}+\frac{12}{x^3}-\frac{100x}1\]That doesn't change its value, of course. Now we multiply each term both top and bottom by all of the denominators that are not present in that term's denominator: \[\frac{x}{(x+10)} * \frac{x^3*1}{x^3*1}+\frac{12}{x^3}*\frac{(x+10)*1}{(x+10)*1}-\frac{100x}{1}*\frac{(x+10)*x^3}{(x+10)*x^3}\] Want to try doing the rest?

OpenStudy (anonymous):

okay, \[x ^{3}+10x ^{2}+12/x ^{3}-100x\]

OpenStudy (anonymous):

is this it ?

OpenStudy (whpalmer4):

Don't think so...how did you get a - sign in the denominator?

OpenStudy (whpalmer4):

Or is that \[x^3+10x^2+\frac{12}{x^3} - 100x\]? Either way, not correct...

OpenStudy (anonymous):

dang it. umm i confused myself

OpenStudy (whpalmer4):

\[\frac{x}{(x+10)} * \frac{x^3*1}{x^3*1}+\frac{12}{x^3}*\frac{(x+10)*1}{(x+10)*1}-\frac{100x}{1}*\frac{(x+10)*x^3}{(x+10)*x^3}\]\[=\frac{x*x^3*1+12*(x+10)*1-100x*(x+10)*x^3}{(x+10)*x^3*1}\]\[=\frac{x^4+12x+120-100x^5-1000x^4}{x^3(x+10)}\]\[=\frac{-100x^5-999x^4+12x+120}{x^3(x+10)}\]

OpenStudy (whpalmer4):

You weren't the only one! :-)